White vinegar is a solution of acetic acid in
water. There are two strengths of white
vinegar- a 5% solution and a 10% solution.
How many milliliters of each solution must
be mixed to make 50 mL of a 9% vinegar
solution?

Answer by stanbon(75887)

There are two strengths of white
vinegar- a 5% solution and a 10% solution.
How many milliliters of each solution must
be mixed to make 50 mL of a 9% vinegar
solution?
------
Equation:
acid + acid = acid
0.05x + 0.10(50-x) = 0.09*50
------
5x + 10*50 - 10x = 9*50
-5x = -50
x = 10 mL (amt. of 5% solution needed)
----
50-x = 40 mL (amt. of 10% solution needed)
==========
Cheers,
Stan H.
===========